OpenStudy (anonymous):

what is the product in simplest form? and state any restrictions. 4a^5/7b^4 * 2b^2/2a^4

OpenStudy (anonymous):

can you help @hartnn

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@JamesJ

OpenStudy (jamesj):

First, \[ 4a^5/7b^4 \times 2 b^2/(2a^4)= 4a^5b^{-4} \times \frac{1}{2} b^2 a^{-4} \] Now collect like terms.

OpenStudy (anonymous):

i don't see any like terms though... the ^-4's

OpenStudy (jamesj):

\[ 4a^5b^{-4} \times \frac{1}{2} b^2 a^{-4} = \left( 4 \times \frac{1}{2} \right) (a^5a^{-4}) (b^{-4}b^2) = ....\]

OpenStudy (jamesj):

and I've dropped a 7 along the way

OpenStudy (jamesj):

\[ \frac{4}{7}a^5b^{-4} \times \frac{1}{2} b^2 a^{-4} = \left( \frac{4}{7} \times \frac{1}{2} \right) (a^5a^{-4}) (b^{-4}b^2) = .... \]

OpenStudy (anonymous):

ok so now what...? i know u would get 4/14 but what about the rest??

OpenStudy (jamesj):

simplify, for example, \[ a^5 a^{-4} = a^{5-4} = a^1 = a \]

OpenStudy (anonymous):

oh wow yeah i see that.... haha so for the b it woudl be b^2?

OpenStudy (jamesj):

-4 + 2 = -2